# ViMRHP: A Vietnamese Benchmark Dataset for Multimodal Review Helpfulness Prediction via Human-AI Collaborative Annotation

### v1

@@ 3.2 Annotation Scheme and Guideline
Helpfulness Score = (Key-aspect + Decision-making advice + Image-helpfulness) / 3 |
@@ Appendix 0.B Instruction
1.0 - Không đề cập đến khía cạnh cụ thể nào của sản phẩm. 2.0 - Đề cập đến một khía cạnh cụ thể của sản phẩm. 3.0 - Đề cập đến hai khía cạnh cụ thể của sản phẩm. 4.0 - Đề cập đến ba khía cạnh cụ thể của sản phẩm. 5.0 - Đề cập đến nhiều hơn bốn khía cạnh cụ thể của sản phẩm. Decision-making advice: Khuyến nghị mua hàng. {decision_making_advice} |
1.0 - Mô tả trải nghiệm cá nhân mơ hồ, không đưa ra khuyến nghị. |
2.0 - Mô tả trải nghiệm cá nhân rõ ràng, nhưng không có khuyến nghị. |
3.0 - Ngầm đưa ra lời khuyên về việc có nên mua sản phẩm hay không. |
4.0 - Đưa ra lời khuyên rõ ràng về quyết định mua hàng. |
5.0 - Đưa ra lời khuyên cụ thể cho từng đối tượng khách hàng. |
@@
1.0 - Does not mention any specific aspect of the product. |
2.0 - Mentions one specific aspect of the product. |
3.0 - Mentions two specific aspects of the product. |
4.0 - Mentions three specific aspects of the product. |
5.0 - Mentions more than four specific aspects of the product. |
@@
1.0 - Describes an ambiguous experience without giving any purchase advice. |
2.0 - Clearly describes the experience but does not provide purchase advice. |
3.0 - Implicitly suggests whether the product is worth buying. |
4.0 - Strongly implies whether the product is worth buying. |
5.0 - Clearly recommends the product, specifying the target users or suitable situations. |
@@
1.0 - Does not meet any criteria. |
2.0 - Meets one criterion. |
3.0 - Meets two criteria. |
4.0 - Meets three criteria. |
5.0 - Meets all four criteria. |

### v2

@@ 3.2 Annotation Scheme and Guideline
| Helpfulness Score = (Key-aspect + Decision-making advice + Image-helpfulness) / 3 | |
@@ Appendix 0.B Instruction
| 1.0 - Không đề cập đến khía cạnh cụ thể nào của sản phẩm. 2.0 - Đề cập đến một khía cạnh cụ thể của sản phẩm. 3.0 - Đề cập đến hai khía cạnh cụ thể của sản phẩm. 4.0 - Đề cập đến ba khía cạnh cụ thể của sản phẩm. 5.0 - Đề cập đến nhiều hơn bốn khía cạnh cụ thể của sản phẩm. Decision-making advice: Khuyến nghị mua hàng. {decision_making_advice} |
| 1.0 - Mô tả trải nghiệm cá nhân mơ hồ, không đưa ra khuyến nghị. |
| 2.0 - Mô tả trải nghiệm cá nhân rõ ràng, nhưng không có khuyến nghị. |
| 3.0 - Ngầm đưa ra lời khuyên về việc có nên mua sản phẩm hay không. |
| 4.0 - Đưa ra lời khuyên rõ ràng về quyết định mua hàng. |
| 5.0 - Đưa ra lời khuyên cụ thể cho từng đối tượng khách hàng. |
@@
| 1.0 - Does not mention any specific aspect of the product. |
| 2.0 - Mentions one specific aspect of the product. |
| 3.0 - Mentions two specific aspects of the product. |
| 4.0 - Mentions three specific aspects of the product. |
| 5.0 - Mentions more than four specific aspects of the product. |
@@
| 1.0 - Describes an ambiguous experience without giving any purchase advice. |
| 2.0 - Clearly describes the experience but does not provide purchase advice. |
| 3.0 - Implicitly suggests whether the product is worth buying. |
| 4.0 - Strongly implies whether the product is worth buying. |
| 5.0 - Clearly recommends the product, specifying the target users or suitable situations. |
@@
| 1.0 - Does not meet any criteria. |
| 2.0 - Meets one criterion. |
| 3.0 - Meets two criteria. |
| 4.0 - Meets three criteria. |
| 5.0 - Meets all four criteria. |